OpenStudy (anonymous):

The product of three and a squared number is twice the sum of the number and four.

OpenStudy (anonymous):

So basically you set this up as: \[3*n^2 = 2(n+4)\] Then solve for n: \[3*n^2 = 2(n+4)\]\[3n^2 = 2n+8\]\[3n^2 - 2n -8 =0\]\[3n^2 - 6n + 4n - 8 = 0\]\[3n(n-2) + 4(n-2) = 0\]\[(3n+4)(n-2)=0\] Then you know that: \[3n+4 = 0\]So: \[n = -4/3\] AND \[n-2=0\]So: \[n=2\] Therefore, n can equal either -4/3 or 2.